Simplify and solve

3. 30 square root of -25 * 8 square root of -49

4. 6 square root of -64 + 12 square root of -36

5. 4 square root of -13 - 6 square root of -13
1
Expert's answer
2013-01-16T10:59:36-0500
We assume that i=sqrt(-1). Knowing that we can solve our task as follows:

3. 30 sqrt(-25) * 8 sqrt(-49)=30*8*sqrt( 25*(-1) )*sqrt(49*(-1) )=30*8*sqrt( 25) *sqrt(-1)*sqrt(49)*sqrt(-1)=240*5*7*i^2=-8400

4. 6 sqrt(-64) +12 sqrt(-36)=6 sqrt(64)*sqrt(-1) + 12 sqrt(36)*sqrt(-1)=( 6 sqrt(64) + 12
sqrt(36) ) *sqrt(-1)=(6*8+12*6)sqrt(-1)=120sqrt(-1)=120i

5. 4 sqrt( -13) -6 sqrt(-13)=(4-6)*sqrt(-13)=-2sqrt(-13)=- 2 sqrt(13)*sqrt(-1)=-2sqrt(13) *i